Grade upto college level Physical Chemistry

Calculate the root mean square velocity of ozone kept in a closed vessel at 20oC and 82 cm mercury pressure.

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
C √3PV/M , P1V1/T1 = P2V2/T2
Where, P1 = 82 cm = 82 * 13.61 * 981 dynes
V1 = Volume of gas at 82 cm and 20oC
P2, V2 and T2 are the conditions at NTP,
=>V1 = P2 V2/T2 * T1/P1 = 76.0 *22400/273 * 293/82 = 22282 cc
M = molecular weight of ozone (O3) = 48
∴ C = √3 *(82 * 13.6 * 981.) * 22282/48
= 3.9 *104 cm sec-1
ALTERNATESOLUTION
Urms = √3RT/M
Given T = 20oC = 20 + 273 = 293K
R = 8.314 * 107 erg per degree per mol
M (of O3) = 48
∴ Urms = √3 * 8.314 * 107 *293/48 = 3.9 * 104 cm sec-1